OpenStudy (anonymous):

A rectangular yard measuring 37ft by 42ft is bordered (and surrounded) by a fence. Inside, a walk that is 2ft wide goes all the way along the fence. Find the area of this walk.

OpenStudy (anonymous):

Well, think about the formula for finding area

OpenStudy (anonymous):

152

OpenStudy (anonymous):

Lenght x width 37x42 =1554

OpenStudy (anonymous):

\[ Area Formulas (Math | Geometry | Area Formulas) (\pi = \pi = 3.141592...) Area Formulas Note: "ab" means "a" multiplied by "b". "a2" means "a squared", which is the same as "a" \times "a". Be careful!! Units count. Use the same units for all measurements. Examples \square = a 2 rectangle = ab parallelogram = bh trapezoid = h/2 (b1 + b2) circle = \pi r 2 ellipse = \pi r1 r2 \]

OpenStudy (anonymous):

39x44=1776 subtract 1554 from 1776 and u get 152

OpenStudy (anonymous):

i hope it rite

OpenStudy (anonymous):

:( aww man :P

OpenStudy (anonymous):

Try watching this video and see if it helps :) :www.khanacademy.org/...area.../perimeter-and-area-basics

OpenStudy (anonymous):

Not the same thing, but it should give an idea on how to solve.

OpenStudy (anonymous):

2*(42+2)+2*(33*2) = 220 220+(33*4) = 132 220+132 = 352 square feet